### v2

# Compositional disorder in a multicomponent non-reciprocal mixture: stability and patterns

Abstract
The mean compositions of individual components can be tuned to control phase behavior in number-conserving passive mixtures. In this work, we investigate the role of variable average density in a system of infinitely many non-reciprocally interacting scalar densities, within the framework of the multi-species non-reciprocal Cahn-Hilliard (NRCH) model. Rather than focusing on specific parameter choices, we study ensembles of systems where the inter-species interaction coefficients and average densities are sampled from probability distributions. We show that non-reciprocity stabilizes the homogeneous mixed state even in the presence of compositional disorder, when compared to the corresponding reciprocal system. Using random matrix theory, we derive a general condition for the onset of spinodal instability, which we verify through simulations. Finally, we illustrate the connection between the statistics of the most unstable eigenvalue and the emergent nonlinear dynamics.
I Introduction
An important question that has driven intense research is – how does a cell achieve self-organization and regulation? A widely accepted route, applicable in many scenarios, is that components in the cell cytoplasm organize into compartments enriched in a specific group of chemicals simply through thermodynamic phase separation [1]. In the past decade, liquid-liquid phase separation has been implicated in the formation of membrane-less organelles [2, 3, 4, 5], regulation of stable structures such as the nucleolus [6, 7, 8], and adaptation to adverse external stimuli in stress granules [9, 10]. It is striking that this intuitively appealing idea surfaced for the first time at the end of the nineteenth century [11, 12] and again in the context of proto-cells [13]. The identification of phase separation as the dominant mechanism opens up new possibilities for regulating condensate formation through pathways accessible in passive systems. These include controlling pH [14], modulating chemical activity [15], coupling to an active fluid [16], and, most notably, altering the mean density of the constituents [17]. The latter presents a particularly promising strategy for regulating phase separation during droplet ripening in the early development of the C. elegans embryo [18].
Figure 1: Role of compositional disorder in multi-species NRCH (a) The average composition of each species is fixed to a randomly chosen value. The dynamical matrix determining the linear stability of the homogeneous mixed state is the sum of a random Gaussian matrix (top) and a diagonal matrix whose entries are determined by the mean density of each species (bottom). In (a) the diagonal entries are drawn from uniform distribution and each color shows a different random value. (b) Three other distributions are chosen to illustrate our main results - the associated probability distribution functions and their name is mentioned. (c) To illustrate the effect of compositional disorder on the dynamics, we run simulations in one dimensional space keeping $\mathbb{M}$ the same in all three cases. The top and the bottom evolve to a state of spatiotemporal chaos in the steady state, while the middle evolves to a phase separated state.
A description of processes within the cell cytoplasm has to incorporate two features: multiple participants and activity in a suitable form. It is well known that the cell interior is a heterogeneous multicomponent environment [19], a feature reflected in numerous studies on passive phase  [20, 21, 22, 23, 24]. However, the perspective that activity must be included in a system-specific form in the formation of bio-organelles is relatively recent [25]. In the field of active matter, active phase separation has been explored through several minimal models that typically involve either a single component—used to model motility-induced phase separation or coupling to chemical reactions [26, 27, 28, 29]—or mixtures with two or more non-reciprocally interacting components [30, 31].
With this background, we take a minimal approach where we model the cell as a multicomponent mixture with non-reciprocal inter-species interactions [32, 33, 34, 35, 36, 37, 38]. Non-reciprocity appears naturally in mixtures of active particles systems: whether polar [39] or scalar [30, 31], and also in quorum sensing mixtures [40, 41], or mixtures of Janus colloids [42]. The hallmark of non-reciprocal number conserving systems is that it combines features of pattern-forming systems [43, 44] and phase separating systems [45, 46]. Moreover, the dynamics can be tuned by the changing composition, thus transitioning from stable mixtures to traveling lattices through a conserved version of the Hopf-bifurcation [30].
We address a question that arises naturally - how does assigning a varying mean density to each species, hereafter referred to as compositional disorder, influence the stability and pattern formation within the system? In [36], it was shown that non-reciprocal activity in the form of asymmetric correlation between the interaction coefficients between the species stabilizes the homogeneous mixed state in comparison with the system with passive interactions. It means that a temperature like control parameter has to be lowered more to drive the active system to a patterned state.
In this work we use random matrix theory [47, 48, 49, 50] to study the linear stability of the system as pioneered in the seminal works of Wigner [51], and May [52]. This approach has been successfully applied to a wide range of physical systems, including the stability of ecosystems with random interactions [53, 54, 55, 56, 57], and disordered systems such as spin glasses [58, 59, 60]. We adopt a statistical perspective, replacing the complexity of individual interaction networks with probabilistically sampled ensembles of systems [21, 61, 62]. Within this framework, the mean composition is treated as a stochastic variable drawn from a distribution, as recently investigated in passive systems [63]. The linearized dynamics of the multi-species NRCH model reduces to studying the stability of the asymmetric matrices with diagonal disorder as represented in Fig. 1(a) facilitating the statistical estimation of the onset of the spinodal instability [64], and the possible phases in this system.
We show that for any choice of compositional disorder, the non-reciprocally interacting system is more stable than the passive one. The theoretical predictions are supported by the numerical solution of the nonlinear model for some choices of the distribution for the average densities shown in Fig. 1(b). We demonstrate the role of complex eigenvalues in producing diverse dynamics that are not permitted in the equilibrium counterpart. The paper is organized as follows: we introduce the theoretical framework of the multi-species NRCH model in Section II. In Section III, we discuss some known results to determine the linear stability of a system with compositional disorder in which case it is essential to consider random Gaussian matrices whose diagonal matrices have been drawn from a different distribution. In the next section, we present some general properties of the dynamical matrix leading to the results in section III. We discuss the varied dynamics seen in the model in section V and end with a summary of the main results in the paper in VI.
II multi-species NRCH with compositional disorder
We adopt the framework presented in [36] for a multicomponent active mixture and consider $N$ individually conserved active species. These species interact via pairwise non-reciprocal interactions originating from the active nature of the system, and reciprocal interactions driven by a free energy. The scalar field $\phi_{a}({\bm{r}},t)$ represents the number density of the $a$-th species. The number conserving densities obey gradient dynamics of the form $\partial_{t}\phi_{a}=\Gamma_{a}\nabla^{2}\mu_{a}$ with mobility $\Gamma_{a}$ and active chemical potentials $\mu_{a}$. $\mu_{a}$ consists of two parts: the first is derivable from a free energy, and the second represents the non-equilibrium active contribution. The passive part is derived from the free energy functional $F=\int\mbox{d}^{d}\bm{r}f$, where $f$ is the free energy density which promotes macroscopic phase separation. We use a minimal polynomial form for $f$ as shown below
$f$ $=$ $\sum_{a}\left[\frac{1}{2}\Theta_{a}\phi_{a}^{2}+\frac{s_{a}}{4}\phi_{a}^{3}-\frac{K}{2}|\bm{\nabla}\phi_{a}|^{2}\right]$ (1)
$\frac{1}{2}\sum_{ab}\chi_{ab}\phi_{a}\phi_{b}.$
$\chi_{ab}$ is Flory-Huggins type interaction parameter between two species $a$ and $b$. To make progress, we have made some assumptions to make the problem more tractable while still retaining enough complexity. The interfacial tension is assumed to be identical for all species and equal to $K$ to focus on effects that are distinct from a Turing type instability that appears when the tensions are chosen to be different [65]. The coefficient of self-interaction $\Theta_{a}=(1+\Theta)$ is the same for all species, akin to an external tuning parameter like the temperature of the system that controls the transition to a patterned state. We now rescale time, space, and the scalar fields as $t\to{K}/{\Gamma}t$, $x\to\sqrt{K}x$, and $\phi_{a}\to\phi_{a}/\sqrt{s_{a}}$, to recast the equations as follows
$\partial_{t}\phi_{a}=\nabla^{2}\left[(1+\Theta)\phi_{a}+\phi_{a}^{3}+\sum_{b}M_{ab}\phi_{b}\right]$
$-\nabla^{4}\phi_{a},\,\,\int\mbox{d}^{d}\bm{r}\phi_{a}(\bm{r},t)=\phi_{a0}.$ (2)
We consider a matrix $\mathbb{M}$ of general interactions whose coefficients are drawn from a Gaussian distribution of zero mean and variance $\sigma$.
$M_{ab}(\chi,\alpha)=\frac{\bar{\alpha}}{\sqrt{2N(\bar{\alpha}^{2}+\bar{\chi}^{2})}}A_{ab}+\frac{\bar{\chi}}{\sqrt{2N(\bar{\alpha}^{2}+\bar{\chi}^{2})}}S_{ab},$ (3)
where $\mathbb{A}$ and $\mathbb{S}$ are fully asymmetric and symmetric matrices respectively, i.e. $A_{ab}=-A_{ba}$, and $S_{ab}=S_{ba}$. Eq. (3) is written such that the variance of the coefficients is $\sigma^{2}$ and the correlation between the diagonal partners determines the level of non-reciprocity in the system, which we call $\alpha$. The properties of $\mathbb{M}$ are summarised as follows
$\langle M_{ab}M_{ab}\rangle=\frac{\sigma^{2}}{2N},\,\,M_{aa}=0,$
$\langle M_{ab}M_{ba}\rangle=\frac{\sigma^{2}}{2N}(1-2\alpha^{2}),\,\,\alpha=\frac{\bar{\alpha}}{\sqrt{\bar{\alpha}^{2}+1}}.$ (4)
where the angular brackets denote averaging over ensembles of $\mathbb{M}$.
To determine the linear stability of the system we substitute $\phi_{a}=\phi_{a0}+\delta\phi_{a}$ in Eqs. (II) and to obtain the linearised dynamics around the homogeneous state
$\partial_{t}\delta\phi_{a}=-q^{2}\sum_{b}\left[M_{ab}+(1+\Theta+3\phi^{2}_{a0})\delta_{ab}\right]\delta\phi_{b}.$ (5)
Defining a diagonal matrix $C_{aa}=3\phi^{2}_{a0}$, whose off-diagonal entries are zero, it is clear from Eq. (5) that the linear stability of the system is determined by the eigenvalues of the matrix $\mathbb{D}=\mathbb{M}+\mathbb{C}$. The effect of varied composition thus enters the dynamical matrix as diagonal disorder through the diagonal matrix $\mathbb{C}$ whose diagonal elements $\bar{\phi}$ are drawn from a distribution $P(\bar{\phi})$. The boundary of the spinodal instability is defined as a part of the multi-dimensional parameter space where at least one of the N eigenvalues is unstable. The onset of the spinodal instability is thus determined by the eigenvalue $\lambda_{0}$ of $\mathbb{D}$ with the smallest real part. With this motivation, we will first discuss the onset of pattern formation in the system.
Figure 2: Spinodal instability in theory and simulations: (a) The boundary of linear stability is plotted with a solid line in the $\Theta-\alpha$ plane, for a fixed set of parameters for the discrete bimodal distribution, see Eq. (13). The predictions are tested using simulations with $50$ realisations of $\mathbb{M}$. The colour of the markers reflects the ensemble and space-averaged order parameter $\Delta$ [see Eq. (9)], which vanishes outside the spinodal shown by a light hue and changes discontinuously to a finite value inside the unstable region, as marked by a darker hue. The grey line is the same for identical mean composition of the densities. The plots show that the relative stability depends both on the distribution and on $\alpha$ (b) Same as panel (a) for the exponential distribution.
III Spinodal
The most unstable eigenvalue $\lambda_{0}$ determines the threshold value of self-interaction or temperature $\Theta_{\rm sp}$ of the linearly stable system,
$\Theta_{\text{sp }}\left(\alpha,\phi_{0},\{D\}\right)=-1-\operatorname{Re}\left[\lambda_{0}\left(\alpha,\{D\}\right)\right],$ (6)
where $\{D\}$ denotes the set of parameters that characterize the distribution $P$. The statistics of $\lambda_{0}$ are determined by the properties of the dynamical matrix $\mathbb{D}$. For $\Theta$ lower than $\Theta_{\rm sp}$, the homogeneous state is unstable and patterns are formed in the steady state.
For $P(\bar{\phi})=\delta(\bar{\phi}-3\phi_{0}^{2})$, i.e. if the mean density of each species is identical, $\lambda_{0}$ is given by the eigenvalues of $\mathbb{M}$ shifted by a constant factor [36].
$\Theta_{\rm sp}=-3\phi_{0}^{2}-\alpha^{2}.$ (7)
Eq. (7) implies that for a fixed $\phi_{0}$, $\Theta_{\rm sp}$ is smaller for the active system in comparison to the passive system ($\alpha=0$). The control parameter $\Theta+1$ is synonymous with the strength of self-interaction. If all entries of the matrix $\mathbb{M}$ were set to zero, the non-interacting system would not phase separate for $\Theta>-1$. $\Theta<-1$ implies attractive interactions between particles of the similar type leading to phase separation. Eq. (6) shows that random passive interactions generate stronger effective attractive interactions between the species, such that to tune the system to a homogeneous state, it is necessary to increase intra-species repulsive interactions. For a fully non-reciprocal system, $\alpha=1$, $\Theta_{\rm sp}$ decreases to $-1$ which is the threshold for the non-interacting system. The non-reciprocal counterpart is thus more stable than the reciprocal system. The factor of unity in Eq. (6) simply sets $\Theta_{\rm sp}$ for passive interactions to zero which serves as a good reference point for comparisons.
We will now state the first important result of our work before illustrating its significance for specific choices of $P$. We find that that for any choice of $P$, $\Theta_{\rm sp}$ satisfies the following simple relation
$\Theta_{\rm sp}=-1-\mbox{Re}[\lambda_{0}]=l_{0}+m\alpha^{2}.$ (8)
The slope $m$ and $l_{0}$ depend only on the parameters $\{D\}$ characterizing the distribution $P$. $l_{0}$ determines the stability of the passive system, $\alpha=0$, in the presence of compositional disorder. To compare the stability of the state with compositional disorder, with $l_{0}>(-3\phi_{0}^{2}+1)$, compositional disorder stabilizes the homogeneously mixed state. The sign of the ‘slope’ $m$ between $\Theta_{\rm sp}$ and $\alpha^{2}$ determines the stability of the active system. For $m<0$ the active system is more stable than the passive one.
In Fig. 2 we have plotted the theoretical prediction for $\Theta_{\rm sp}$ for the bimodal and exponential distributions described in IV, see Eqs. (13) and (14). To test the predictions of linear stability analysis, we solve Eq. (II) numerically. We start our simulations with uniform composition $\phi_{a0}$ such that $3\phi_{a0}^{2}$ is drawn from either the bimodal or the exponential distributions. The simulations are initiated with random Gaussian perturbations. For other details on system size, methods used, see the Appendix IV. The results of the simulation are shown in Fig. 2, where the markers denote the points where the simulations are carried out and the colour of the markers carries information about the magnitude of the order parameter.
$\Delta=\int\mbox{d}x\frac{1}{N}\sum_{a}\left(\phi_{a}(x,t)-\phi_{a0}\right)^{2}.$ (9)
We run $50$ realizations of our simulation for $N=100$ components to obtain the results in Fig. 2.
Figure 3: Boundary enclosing the eigenvalues in the complex plane (a) The boundary and the spectrum of a large random matrix with diagonal disorder drawn from the bimodal distribution, see Eq. (13) for $d_{2}=-d_{1}=d$. (b) The diagonal elements of $\mathbb{C}$ are drawn from a uniform distribution where $d_{-}=1$ and $d_{+}=2$. (c) The diagonal disorder is drawn from an exponential distribution with $d=0.6$. (d) The spectrum and the boundary of support for a large random matrix with diagonal elements from a beta distribution with $a=2$ and $b=5$. For all examples, the random matrix is of the size $\text{N}=1000$ and $\chi=1,\alpha=0.5$. (e) The diagonal disorder are drawn from a trimodal distribution with $d_{1}=0.5$, $d_{2}=2$, $d_{3}=5$, $p_{1}=\frac{1}{2}$ and $p_{2}=\frac{1}{5}$ the random matrix is of a size $N=2000$ and $\chi=1$.
Figure 4: The slope $m$ and the intercept $l_{0}$ for different types of compositional disorder. The slope $m$, referred to in Eq. (8) and (IV) for the definition, is calculated for the distributions indicated in the figure. The heatmap of $m$ for the two parameters in (a) and (d) and its variation with the width $d$ in (b) and (c) shows that it is always less than zero. Both the uniform and exponential distributions approach the delta function for $d\to 0$ so that $m$ approaches $-1$ in panels (b) and (c). Panels (e) and (f) show that $l_{0}$ can be positive or negative, meaning that $\Theta_{\rm sp}$ for the passive system can be higher or lower for the corresponding system with identical mean densities.
IV Eigenvalue spectrum
In this section, we will explore four different distributions which are either bounded or unbounded, and which assume either discrete or continuous values to further explore the results in the previous section. The boundary enclosing the spectrum of complex eigenvalues is well approximated by a continuous curve when $N\to\infty$, see Fig. 3 for some examples. In Fig. 3, the markers denote eigenvalues of a single realization of the matrix $\mathbb{D}$ while the boundary has been calculated using methods described in [66], which we will briefly describe here. The spectral distribution $\rho$ of the eigenvalues $\lambda_{j}$ of $\mathbb{D}$ can be written formally as
$\rho(z)=\lim_{n\rightarrow\infty}\frac{1}{n}\left\langle\sum_{j=1}^{n}\delta\left(x-\operatorname{Re}\left(\lambda_{j}\right)\right)\delta\left(y-\operatorname{Im}\left(\lambda_{j}\right)\right)\right\rangle,$ (10)
The boundary of the spectrum where $\rho(z)$ is be called $S$ and is given by the points $z\equiv(x,y)$ in the complex plane. The starting point of our analysis is the coupled equations determining the contour of points $(x,y)$ in the complex plane that envelopes the spectrum of $\mathbb{D}$. To do so, it is essential to introduce another quantity $\mbox{Re}(g)$ the physical meaning of which has been mentioned in the Appendix VII.
$\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\bar{\phi}-x+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}}{\left((\bar{\phi}-x)+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}\right)^{2}+{y^{2}}/{(2\alpha)^{2}}}$
$=-\operatorname{Re}\left(g\right),$ (11)
and
$\int\mbox{d}\bar{\phi}P(\bar{\phi})\,\frac{\sigma^{2}}{\left((\bar{\phi}-x)+\operatorname{Re}\left(g\right)(1-2\alpha^{2})\sigma^{2}\right)^{2}+{y^{2}}/{(2\alpha)^{2}}}$
$=1.$ (12)
For $P(\bar{\phi})=\delta(\bar{\phi}-\phi_{0})$, the eigenvalues are distributed uniformly with an ellipse with major and minor axes equal to $1-\alpha^{2}$ and $\alpha^{2}$ respectively, such that $\mbox{Re}[\lambda_{0}]=1-\alpha^{2}$ as can be verified readily from Eqs. (11) and (12).
As the name suggests, the discrete bimodal distribution is a discrete distribution in a bounded domain with delta peaks at two values $d_{1}$ and $d_{2}$.
$P(\bar{\phi})=p\delta(\bar{\phi}-d_{1})+(1-p)\delta(\bar{\phi}-d_{2}).$ (13)
As seen in Fig. 3 (a), the bimodal effectively splits the system into two groups containing $pN$ and $(1-p)N$ species each. For a fixed $p$ and $d$ close to 00 we find that the boundary of the spectrum simply elongates in the direction of the real axis. As $d$ is increased further, a symmetric fissure appears in the boundary before it is symmetrically split into two contours. The numerical results match well with the contour determined from solving Eqs. (11) and (12).
The exponential distribution is unbounded and associated with a single parameter $d$ that determines the upper-limit of the values that are sampled with comparatively higher probability.
$P(\bar{\phi})=\frac{1}{d}\exp\left(-\frac{\bar{\phi}}{d}\right),\,\,d>0.$ (14)
The boundary in Fig. 3 (b) reflects the asymmetry of the distribution and shows a tail that reflects the large deviations due to a few species with high mean composition.
The uniform distribution is defined by its width $d$ within which all compositions are sampled with the same probability
$P(\bar{\phi})=\frac{1}{d_{2}-d_{1}},\,\,\bar{\phi}\in[d_{1},d_{2}],$ (15)
shown in Fig. 3 (c) shows an elongated boundary along the real axis. With increasing $d$, the aspect ratio increases as well.
Lastly, we examine the Beta distribution, which is a continuous and bounded distribution parameterized by indices $a$, $b$, and its bound $d$.
$P(\bar{\phi})=N_{a,b}^{-1}\ \bar{\phi}^{a}\left(d-\bar{\phi}\right)^{b},\quad x\in\left[0,d\right]$
$N_{a,b}=d^{a+b+1}B(a+1,b+1),$ (16)
where $N_{a,b}$ is the normalization constant and $B$ is the Beta function.
We next motivate the general relation in Eq. (8) by a closer examination of the structure of the coupled equations Eqs. (11) and (12) and verify it by numerical calculations using the distributions introduced above. To do so, we define new variables
$\tilde{x}=x-\mbox{Re}(g)(1-2\alpha^{2})\sigma^{2},\,\,\tilde{y}=\frac{y}{2\alpha^{2}},$ (17)
using which we can re-write Eq. (11) as
$\mbox{Re}(g)=-\int\mbox{d}\bar{\phi}\,P(\bar{\phi})\frac{(\bar{\phi}-\tilde{x})}{(\bar{\phi}-\tilde{x})^{2}+\tilde{y}^{2}}\equiv f_{\tilde{y}}(\tilde{x}),$ (18)
where the equivalence defines the function $f_{\tilde{y}}$ whose form depends on $\tilde{y}$. Similarly, Eq. (12) can be re-expressed using a function $g_{\tilde{y}}$
$\frac{1}{\sigma^{2}}=\int\mbox{d}\bar{\phi}\,P(\bar{\phi})\frac{1}{(\bar{\phi}-\tilde{x})^{2}+\tilde{y}^{2}}\equiv g_{\tilde{y}}(\tilde{x}),$ (19)
where the subscript $y$ denotes the value of $\tilde{y}$ at which the integral Eq. (19) is evaluated. Eq. (19) can be inverted to obtain
$\tilde{x}$ $=$ $g_{\tilde{y}}\left(-\sigma^{-2}\right),$ (20)
As seen in Fig. 3 (a), $y$ Similarly inverting Eq. (18), we have $\mbox{Re}(g)=f^{-1}_{\tilde{y}}(\tilde{x})$, which can be used together with Eq. (20) to obtain
$\mbox{Re}(g)$ $=$ $f^{-1}_{\tilde{y}}\left[g_{\tilde{y}}\left(-\sigma^{-2}\right)\right].$ (21)
for a given $\tilde{y}=y/\alpha^{2}$, we can in principle determine the functions $f_{\tilde{y}}$ and $g_{\tilde{y}}$ which can be used to determine $\tilde{x}$ and $\mbox{Re}(g)$ to obtain $x$ as
$x=g_{\tilde{y}}\left(-\frac{1}{\sigma^{2}}\right)+\mbox{Re}(g)(1-2\alpha^{2})\sigma^{2}.$ (22)
From Eq. (17) we deduce that the curves $S$ for two different values of $\alpha$, the $y$ values are proportional. For $\tilde{y}=0$, i.e., the points at which $S$ intersects the $x$ axis are the same for $S$ for any value of $\alpha$. This means that the topology of $S$ is determined by $\{D\}$ only and does not depend on $\alpha$ at all, as illustrated using the bimodal distribution in panel (e) of Fig. 3.
By the definition of $\lambda_{0}$, we have using Eq. (22)
$\lambda_{0}=L_{0}-m\alpha^{2}.$ (23)
where
$L_{0}=g_{0}\left(-\sigma^{-2}\right)+f^{-1}_{0}\left[g_{0}\left(-\sigma^{-2}\right)\right],$
$m=2\sigma^{2}\mbox{Re}(g)=2\sigma^{2}f^{-1}_{0}\left[g_{0}\left(-\sigma^{-2}\right)\right].$ (24)
We note that $\alpha=1/\sqrt{2}$ is a special point because at this point $\tilde{x}=x$, and it is particularly easy to carry out the algebraic operations described above, as shown with a few examples in the Appendix VII. It is also relatively easy to discuss the sign of $m$ at this value of $\alpha$. The diagonal disorders that we consider are always positive quantities and thus bounded by zero from below. From Eq. (18) it is clear that at the extreme value $\lambda_{0}$ of $\tilde{x}=x$, the sign of the integrand is determined by $(x-\bar{\phi})$, which is always negative in the domain of $\bar{\phi}$.
To test that the predictions are correct, we plot the curvature $m$ by varying the parameters $\{D\}$ for the distribution shown in Fig. 1. For each value indicated in the axes of the panels in Fig. 4, we calculate $\lambda_{0}$ at two values of $\alpha^{2}$ and calculate the curvature and verify that it is indeed positive everywhere.
$m=\frac{\lambda_{01}-\lambda_{02}}{\alpha^{2}_{2}-\alpha^{2}_{1}}.$ (25)
Considering again the point $\alpha=1/\sqrt{2}$, we can argue that the topology of the curve is restrained in a particular way.
$\frac{x(\alpha_{1})-x(\alpha_{2})}{\alpha_{2}^{2}-\alpha_{1}^{2}}=2\mbox{Re}(g)(\tilde{y})\sigma^{2}$ (26)
Eq. (22) suggests that the curve $S$ can be calculated at all values of $\alpha$ if we calculate it at $\alpha=1/\sqrt{2}$. For the bimodal distribution, the curve in $S$ splits into two above a threshold value of $d$, meaning that $\tilde{y}$ vanishes at four points on the $x-$axis. The topology of the boundary is preserved for all values of $\alpha$ as shown in Fig. 3(e) for a trimodal distribution where the $\{D\}$ are fixed and $\alpha$ only is varied. The spectrum for the values indicated at the values of alpha indeed follow the insight from theory.
The points at which $S$ intersects the real axis are unchanged for any value of $\alpha$.
Figure 5: Statistics of $\lambda_{0}.$ In the limiting case of infinite conserved densities, $\lambda_{0}$ is real. For any finite sized system, $\langle\lambda_{0}\rangle$ averaged over ensembles of $\mathbb{D}$, is in fact complex. (a) The imaginary and reals parts, $\mbox{Im}(\lambda_{0})$ and $\mbox{Re}(\lambda_{0})$ are plotted as a function of $\alpha$, encoded in the colour of the marker. The error bar indicates the variance. At each value of $\alpha$, we consider $10^{4}$ realizations of $\mathbb{D}$ for $N=300$. From the plots, it is clear that the variability in $\mbox{Im}(\lambda_{0})$ depends on the type of compositional disorder. Notice also that the variations are increased compared to the uniform average composition. (b) The fraction of complex eigenvalues are plotted as a function of $\alpha$ shows that the fraction increases with $\alpha$ although it does not approach unity.
Figure 6: Pattern formation with decreasing $\Theta$. Panel (a) illustrates the dynamics in the steady state for a fixed randomly sampled realisation of $\mathbb{M}$ and $\mathbb{C}$ for a system with $N=100$ with diagonal chosen from a discrete bimodal distribution. The values chosen for $\Theta$ are indicated in the figure. We generically find spatiotemporal chaos in ten staedy state. (b) illustrates the dynamics in the steady state with the diagonal disorder from an exponential distribution. Each kymograph shows the evolution of the scalar field for one species among $N$ of them.
V Imaginary eigenvalues and steady state dynamics
In this section, we further characterize $\lambda_{0}$ for a finite number of components $N$ and relate the trends that we observe in the statistics of $\lambda_{0}$ to patterns in numerical solutions of the full nonlinear model in Eq. (II). $\lambda_{0}$ is real only when $N\to\infty$ a limit that represents a theoretical abstraction. In reality, the statistics of $\lambda_{0}$ exhibit strong finite size effects that are relevant to real experimental systems and dynamics observed in numerical simulations. The most important aspect of this is that $\lambda_{0}$ is imaginary with a significant probability - an important fact because many novel features of the multi-species NRCH model can be attributed to the emergence of complex eigenvalues in the linearised dynamics of a phase separating mixture.
In Fig. 6(a) the mean value of $\lambda_{0}$ averaged over many realizations of the matrix $\mathbb{M}$ is shown for the exponential, discrete bimodal, and the identical distributions. The variance of $\lambda_{0}$, indicated by the error bar, is larger for the first two in comparison with the identical distribution. The important difference in the average value of $\lambda_{0}$ stems from the fact that in the absence of compositional disorder, $\mathbb{D}$ has only imaginary eigenvalues. For $\alpha\to 1$ we expect that $\mbox{Re}(\lambda_{0})=0$, meaning that all eigenvalues including $\lambda_{0}$ are now strictly complex such that the fraction of complex $\lambda_{0}$ approaches unity as seen in Fig. 6(b). In the other two cases, the diagonal disorder of $\mathbb{D}$ creates the possibility that $\lambda_{0}$ is real, meaning that the probability is capped by a limiting value as seen in Fig. 6(b). For the exponential, we expect that as $d\to 0$ meaning that the range of allowed values of $\bar{\phi}$ shrinks, the curve will bend upward and resemble the case with no compositional disorder.
Finally we show a few examples of the dynamics for a few chosen values of $\Theta$ below $\Theta_{\rm sp}$. We note that compositional disorder plays a role in the steady state dynamics as seen in Fig. 1(c) as seen in the difference between the kymographs representing the space-time dynamics for compositional disorder sampled from the bimodal distributions and the exponential in panels (a) and (b) respectively of Fig. 2. The same interaction matrix $\mathbb{M}$ has been chosen for the two panels (a) and (b) in Fig. 2. The parameters for the two choices of disorder are mention in the caption of the Fig. 2. The self-interaction $\Theta$ is lowered (see values indicated in the panels) to showcase the dependence of pattern on the difference $\Theta_{\rm sp}-\Theta$. We generically find spatiotemporal chaos in the steady state.
VI Conclusions
To conclude, we have explored the role of compositional disorder in determining the onset of the spinodal instability and pattern formation in a random non-reciprocally interacting mixture. Using tools of random matrix theory, we have shown that the instability threshold goes down quadratically with the strength of non-reciprocity in a system with infinitely many species. The stability of a system with any compositional disorder is determined by the interplay of two parameters - the coefficient multiplying the square of the activity, and the threshold value of the effective temperature of the same system at zero activity. These parameters are determined solely by the distribution from which the disorder is drawn from. Through analytical arguments, we have motivated the dependence of this particular form and provided arguments for the relative stability of the active system compared to its passive counterpart. These predictions are validated through extensive numerical simulations of two types: (1) generating ensembles of the dynamical matrix to analyze the statistical properties of the most unstable eigenvalue, and (2) performing direct numerical simulations of the multi-species NRCH equation.
Previous work has shown that non-reciprocity plays a key role in tuning the system to the threshold of instability [36]. However, it is reasonable to assume that controlling the volume fractions of chemicals in a mixture—and thereby regulating phase separation—may be a more accessible experimental pathway. Our results, which demonstrate diverse dynamical behaviors arising from simple changes in the form of compositional disorder, are promising. This work represents a first step toward understanding the composition dependence of multi-component active mixtures. A detailed exploration of the resulting dynamical phases will be presented in future studies.
Acknowledgements.
LP and SS acknowledge useful discussions and support from Ramin Golestanian and Navdeep Rana.
VII Appendix
In this Appendix, we will describe the methods used to obtain Eq. (8) of the main text. The resolvent for the dynamical matrix $\mathbb{D}$ is defined as
$G(z)=\int\mathrm{\leavevmode\nobreak\ d}z^{\prime}\frac{\rho\left(z^{\prime}\right)}{z^{\prime}-z}$ (27)
where $\rho(z)$ is the spectral density of $\mathbb{D}$ and the integral is carried out in the complex plane. As the resolvent is singular whenever $z$ is equal to one of the eigenvalues, we have to regularize the resolvent by considering the quaternionic resolvent $G(q)$, as a function of the quaternionic argument $q$ [67, 68].
$G(q)=\int\mathrm{\leavevmode\nobreak\ d}z^{\prime}\rho\left(z^{\prime}\right)\left(z^{\prime}-q\right)^{-1}$ (28)
which is well-defined as long as $q=z+uj$ with $u\neq 0$. Using the generalization of Pastur’s relation [69] for the asymmetrical matrices which was studied in [67] and the properties of the quaternions, we can rewrite the resolvent as
$g+\beta\mathrm{j}=\int\mathrm{\leavevmode\nobreak\ d}\bar{\phi}P(\bar{\phi})\frac{(\bar{\phi}-{z^{*}}-(1-2\alpha^{2})g^{*})+\beta\mathrm{j}}{|\bar{\phi}-z-(1-2\alpha^{2})g|^{2}+|\beta|^{2}},$ (29)
where both sides of this equation are quaternionic. Writing the complex number $g$ as $g=\text{Re}(g)+\text{Im}(g)$ distinguishing the real and the imaginary parts, and $z=x+yi$, we can distinguish the real and imaginary parts of Eq. (29) to obtain the coupled equations for $z$ and $\mbox{Re}(g)$, see Eqs. (11) and (12), in the main text. In arriving at Eqs. (11) and (12) we have further assumed that the diagonal entries of $\mathbb{C}$, and thus the values $\bar{\phi}$ assumes are real.
VII.1 Numerical schemes and parameter choices
In this section, we will describe the methods used to solve the $N$ nonlinearly coupled PDEs to verify the results of the linear stability analysis and to garner some idea of the rich pattern forming ability of these systems of equations.
The simulations shown in Fig. 1 and used in Fig. 4 have been performed using a pseudo-spectral method in a one-dimensional system on a domain of length L discretized with nx grid points to obtain a stable solution of the nonlinear partial differential equations for the N-species equations of motion. For time discretizing, we use the second-order Exponential Time Differencing (ETD2) scheme [70] with a fixed time step $\Delta t$, a small enough time step to ensure reaching a steady state.
For the results in the Fig. (2), the simulation was carried out for $N=100$ components. The system size is chosen to be $L=4096$ and $nx=4096$ in the simulations. The time step for simulations reported in this figure is $\Delta t=2.5\times 10^{-3}$.
The results of the Fig. (6), is for a simulation of $N=100$ components with the system size $L=512$ and $nx=512$ in the simulations. The time step for simulations reported in this figure is $\Delta t=5\times 10^{-3}$. For both results in Fig (2) and (6) the surface tension parameter $K=0.1$.
In all simulations for the discrete bimodal distribution in we chose $d_{1}=0.2$ and $d_{2}=1$. Accordingly, for the exponential distribution that was used in the simulation, we chose $d=0.6$.
In Fig. (6), for the discrete bimodal distribution, from the linear stability analysis we can find $\Theta_{sp}=0.06$ and for the exponential distribution is $\Theta_{sp}=-0.19$.
VII.2 Discrete Bimodal distribution
As a first example, we sample the diagonal entries of $\mathbb{C}$ from the bimodal distribution.
$P(\bar{\phi})=p\delta\left(\bar{\phi}-d_{1}\right)+(1-p)\delta\left(\bar{\phi}-d_{2}\right).$
We assume $d_{1}<d_{2}$ and $d_{1}>0$. Eqs. 11 and 12 assume simple forms for this choice, such that $\lambda_{0}$ can be found analytically setting $y=0$, and using the re-defined variable $\tilde{x}$ we get
$\begin{gathered}1=\frac{p}{\left(\tilde{x}-d_{1}\right)^{2}}+\frac{(1-p)}{\left(\tilde{x}-d_{2}\right)^{2}}\\ \operatorname{Re}\left(g\right)=-\frac{p}{\left(\tilde{x}-d_{1}\right)}-\frac{(1-p)}{\left(\tilde{x}-d_{2}\right)}\end{gathered}$ (30)
The smallest eigenvalue is then obtained as $\lambda_{0}=\tilde{x}-d_{2}-(1-2\alpha^{2})\sigma^{2}\text{Re}(g)$. However, it is worth mentioning that the boundary of support plotted in the 3 is for $\alpha=\sqrt{1/2}$ for which the boundary follows a simpler equation
$f(\lambda)=\frac{p}{|\lambda-d_{1}|^{2}}+\frac{(1-p)}{|\lambda-d_{2}|^{2}}=1/\sigma^{2}$ (31)
where $\lambda=\mbox{Re}(\lambda)+i\mbox{Im}(\lambda)$.